# Towards the Automated Extraction and Refactoring of NoSQL Schemas from Application Code

### v2

@@ 1 Introduction
As far as we know, no approaches applying code analysis to infer NoSQL complete schemas have been published. However, static code analysis of MongoDB applications has been used for related purposes, such as discovering entity evolution and detecting database access operations [20, 5]. [20] presented a strategy aimed at tracking how the schema of a specific entity type evolves throughout an application’s lifecycle. Their approach analyzes different versions of Java applications to identify structural changes over time —i.e., entity versions. However, it does not extract a complete database schema, as relationships between entities are not identified. Boris Cherry et al. proposed a method to detect MongoDB access operations in JavaScript applications. Their work focuses on locating access points—such as queries, insertions, updates, and deletions—across diverse codebases.
@@
    To the best of our knowledge, this is the first code analysis approach capable of extracting NoSQL logical schemas, where entity types are connected through both aggregation and reference relationships. Additionally, the approach supports relationship types. By representing schemas using a generic metamodel, our method becomes applicable to multi-model database tools, enabling broader use in diverse database environments.
@@
    The extracted information in the devised reverse engineering process also enables the automation of database refactorings. In this paper, we illustrate this capability with a strategy to duplicate fields in order to eliminate the need for expensive joins, thus improving query performance.
@@
    The proposed reverse engineering process leverages metamodel-based abstractions [3] to represent the involved information: source code, control flow, database operations, and data structures. The metamodels were designed to ensure platform independence and reusability. In particular, the code metamodel captures common elements of object-oriented languages and can be extended with language-specific constructs.
@@
The full implementation of our approach, which includes all metamodels, model-to-model transformations, and code analysis algorithms used for generating intermediate models, extracting schemas, and automating refactorings, is publicly available on GitHub  [^7] U-Schema Inference and Refactoring Repository: https://github.com/modelum/uschema-code-analysis.
@@ 2 Overview of the approach
User objects contain a watchedMovies field, which holds an array of embedded objects. Each embedded object includes three fields: _id, which records the identifier or key; movie_id, which stores the identifier of the watched movie (i.e., a reference); and stars, which contains the user’s rating of that movie.
@@
1. user = db("Users").query(name == "Brian")
2. if (user.watchedMovies[0].stars >= 5)
println user.name + user.surname + user.email
println "Lastwatchedmovie":
println user.watchedMovies[0].movie_title
@@
1const MongoDB = require(’mongodb’).MongoClient;
@@
3const url = ’mongodb://modelum.es/db:27017’;
4const dbName = ’streamingservice’;
@@
6const client = new MongoDB(url);
7client.connect(err => {
@@ 3 Obtaining an abstract representation of the code
In this section, we describe how the source code to be analyzed is internally represented using two complementary metamodels: the Code metamodel, which captures the structural elements of the program, and the Control Flow metamodel, which models the execution order of its statements. These representations provide the foundation for subsequent analysis stages, including the identification of database operations and the extraction of the schema.
@@ 3.1 From source code to Code models
Instead of relying directly on the abstract syntax tree (AST) of the source code, we define the Code metamodel to represent code in a more abstract and uniform way, independent of the concrete syntax. An excerpt of this metamodel is shown in Figure 4, which illustrates the main concepts and relationships common to object-oriented languages, and specifically includes the JavaScript statements relevant to our work. This decision aligns with the model-driven engineering principle of building metamodels to define the abstract syntax of software languages [12, 14, 26]. In designing the Code metamodel, we were primarily inspired by the Java MODISCO metamodel [4], and to a lesser extent by the Code package of KDM [21]. It is important to note that our metamodel was not tailored to any specific object-oriented language. Instead, language-specific elements are defined in separate extension metamodels, allowing for a modular architecture that can be easily adapted to different programming languages. This design also ensures that the most common statements are centralized in the core Code model, promoting consistency and reuse across language variants.
Next, we describe the Code metamodel in sufficient detail to explain how Control Flow models are derived. A Code model represents an executable unit, such as a JavaScript script or program. It aggregates Containers, global variable declarations (VariableDecl), and Types, which can be either primitive types or classes. A Container represents a structural unit that holds scripts or classes, such as packages, namespaces, folders, or files. Containers can be nested and may aggregate CodeContainers, which in turn group code blocks, class declarations, and variable declarations. A CodeBlock contains an ordered list of Statements, such as conditionals or loops, and also declares local variables. Figure 4 shows only the types of statements that appear in the running example: conditional selection, method calls, variable accesses, and object creation expressions. A special kind of code block is the CallableBlock, which represents executable blocks invoked by the program, such as methods, functions, constructors, and lambda expressions.
Code models are obtained (injected) from the AST provided by a parser. Currently, we have implemented an injector for JavaScript, using the Esprima parser [^8] Esprima parser website: https://www.esprima.org. This parser provides the AST as a set of JSON documents. Notably, each JSON node includes a class field that clearly identifies the type of statement or expression it represents, which facilitates the mapping process. As a result, the correspondence between the Esprima AST elements and the elements of the Code metamodel is simple and direct in most cases, requiring minimal transformation logic.
@@
To validate this stage, we applied the testing strategy defined in [9], which describes a model-driven software reengineering process. A set of simple tests was performed on minimal code snippets, each containing only the essential instructions needed to represent a specific code construct (e.g., a loop). In each test, the code was automatically regenerated from the Code model and then compared with the original code, taking formatting into account. This approach enabled an iterative development process, where the injection logic for a single statement type was implemented and tested in each iteration. A text comparison tool, such as the standard ’diff’ utility on Ubuntu, was used to compare the original and regenerated code. It is worth noting that the incremental development of the Code metamodel prevented us from using a model-based language workbench to automatically generate the model injector from an EBNF-like JavaScript grammar specification. This decision was motivated by the complexity and large size of the metamodel.
@@ 3.2 Representing the control flow
Code analysis typically requires not only a representation of the syntax tree, but also knowledge of the control flow graph. To represent control flow in our approach, we have defined the metamodel shown in Figure 6. This metamodel was designed based on the representation proposed in the algorithm described in [1]. A Control Flow model is derived from a Code model, with its nodes and edges referencing the corresponding statements in the Code model. Both models serve as input to the code analysis process described in the following section. It is important to note that this representation is not a model of the program’s runtime behavior, but rather a control flow model—that is, a structural abstraction of the code’s possible execution paths. This distinction is particularly relevant in the case of JavaScript, where asynchronous constructs may cause the actual execution order to diverge from the control structure represented in the model.
@@
Data: codeModel : Code model
Result: cfModel : Control Flow Model
1
[…]
3 $codeBlocks\leftarrow getCodeBlocks(codeModel)$
4
5
6foreach $cBlock\in codeBlocks$ do
7       $sGraph\leftarrow createSubGraph(controlFlow)$
8       $sNode\leftarrow createStartNode(sGraph)$
9       $lastNode\leftarrow sNode$
10       $eNode\leftarrow createEndNode()$
11
12      foreach $st\in cBlock.statements$ do
13             $node\leftarrow createConnectedNode(st,st.variables)$
14             $createEdge(lastNode,node)$
15             $lastNode\leftarrow node$
16
17       end foreach
18
19 end foreach
20
21
22 Function branch($st,lastNode$):
23       foreach $newST\in st.statements$ do
24             $node\leftarrow createNode(newST,newST.variables)$
25             $createEdge(lastNode,node)$
26             $lastNode\leftarrow node$
27
28       end foreach
29
30      $endNode\leftarrow createEndNode()$
31       $createEdge(lastNode,endNode)$
32
33      return endNode
34
35
36
37 Function createConnectedNode($st,variables$):
38       $node\leftarrow createNode(st,variables)$
39
40      if $hasArguments(st)$ then
41             foreach $arg\in st.arguments$ do
42                   $createEdge(node,$ $createNode(arg,arg.variables))$
43
44             end foreach
45
46       else if $isSelection(st)$ then
47             $bNode\leftarrow createStartNode(node)$
48             foreach $case\in st.cases$ do
49                   $nNode\leftarrow branch(case.statements,bNode)$
50                   $createEdge(bNode,nNode)$
51
52             end foreach
53
54       else if $isTry(st)$ then
55             $bNode\leftarrow createStartNode(node)$
56             foreach $catch\in st.catchs$ do
57                   $nNode\leftarrow branch(catch.statements,bNode)$
58                   $createEdge(bNode,nNode)$
59
60             end foreach
61
62       else if $isLoop(st)$ then
63             $nNode\leftarrow branch(loop.statements,node)$
64             $createEdge(nNode,node)$
65
66       end if
67      return node
68
69
70
Algorithm 1 Control Flow Construction Algorithm.
Since the Control Flow model is explicitly linked to the Code model from which it is derived, code analysis can traverse control paths to access corresponding statements and perform higher-level reasoning over the program structure.
To generate Control Flow models, we adapted the algorithm described in [1]. In our case, the input is a Code model rather than an abstract syntax tree (AST), and the output is a corresponding Control Flow model. To this end, we defined Algorithm 1, which traverses the input Code model and creates the corresponding nodes and edges in the output Control Flow model. The specific types of nodes and edges generated depend on the kind of element being processed in the Code model.
Algorithm 1 operates as follows. First, the root element of the ControlFlowModel is created (line 1), and all code blocks (CodeBlocks and CallableBlocks) in the Code model are retrieved and collected into a list that is iterated over (lines 2 and 4). For each code block, a corresponding SubGraph is created and initialized with its start and end nodes (lines 5–8). The statements within each code block are then iterated (line 9), and a node is created for each statement using the createConnectedNode function (line 10), which internally calls createNode() (line 25). Each node maintains a reference to the corresponding code statement and the variables used in it. By default, each newly created node is connected to the previously created node (line 11). However, if the current statement is a conditional, loop, or exception trigger, the createConnectedNode function creates new subgraphs to represent the nested statements before establishing the corresponding control flow edges (lines 26–46). As an example, a conditional statement is handled as follows (lines 31–36): First, the initial decision node is created (line 32); then, a new branch is created for each Case (line 34), and a conditional expression edge is added to hold the condition for each branch (line 35). The Branch function is responsible for instantiating the nodes that belong to a specific branch, that is, the statements contained within the code block associated with that branch (lines 17 to 21), and also for creating the final node of the branch (line 22). Finally, an edge is created to connect the last node in the flow to the graph’s end node (line 23).
@@ 64             $createEdge(nNode,node)$
We validated this second step by visually verifying that the generated models accurately represent the control flow of the code. To facilitate this task, the models were stored in a Neo4J graph database, allowing us to leverage the Neo4J Browser  [^9] Neo4j Browser: https://neo4j.com/developer/neo4j-browser, which displays graph query results as navigable visual graphs. Figure 8 shows an excerpt of the graph corresponding to the Control Flow model of the running example. Each node is labeled with a snippet of code. In particular, the graph illustrates the control flow of the if-then statement, which includes a block of three console.log statements. The if node has two outgoing edges: one of type Selection, representing the condition, and another of type Jump, indicating the continuation point if the condition is not met. Each print-related node includes an outgoing edge labeled NEXT, which represents the sequential execution flow, and an additional edge labeled argument, pointing to a node that represents the argument of the print statement. These Neo4J visualizations are significantly more readable and intuitive than the raw Control Flow model, thereby facilitating the validation process. The mapping from the Control Flow model to the Neo4J database is straightforward, and the generation code is produced automatically using the same mechanism employed during the validation of the Code model injection.
@@ 4 Finding Operations and Structure of the database
Figure 9: Database Operation&Structure Metamodel.
Code and Control Flow models are analyzed to discover the implicit database schema and to apply database refactorings. In the first step of the analysis, information about CRUD operations and the structure of the manipulated data is captured in an intermediate representation. This representation is defined by the Database Operation and Structure (DOS) metamodel, as shown in Figure 9. In this section, we first present the algorithm designed to obtain a DOS model. In the following section, we describe how the DOS model is transformed into a U-Schema model, that is, a unified logical schema. The subsequent section illustrates how the DOS model can also be used to support the join query removal refactoring, aiming at improving query efficiency.
@@ 5 $createReferences(dboNodes)$
8 Function backwardTraverse($dboNodes$):
@@ 14             $sNode\leftarrow getPreviousNode(dboNode)$
15             while $\exists$ $sNode$ do
16                   if $isDatabaseOperation(sNode)$land $getReturnVariable(sNode)\in sList$ then
@@ 20                         $markJoinQuery(pDBO)$
22                   else if $isAssignment(sNode)$land $getLeftVariable(sNode)\in sList$ then
@@ 27                  $sNode\leftarrow getPreviousNode(sNode)$
36 Function forwardTraverse($dboNodes$):
@@ 44             $tNode\leftarrow getNextNode(dboNode)$
45             while $\exists$ $tNode$ do
@@ 52                  $tNode\leftarrow getNextNode(tNode)$
61 Function createReferences($dboNodes$):
@@ 66             $createReference(sField,tField)$
Before traversing the control flow graph, two preliminary operations are performed. First, a DOSmodel root element is created (line 1). Second, each subgraph — representing either a function or a script) is traversed to identify the Call nodes associated with database operations. These nodes are collected in the ordered list dbCallNodes (line 2), which reflects the execution order of the program. Relevant nodes are identified by matching function names against those defined in the database management API used in the code. It is important to note that in a Control Flow model, each subgraph is connected to the next in the execution sequence through a CallEdge, as shown in Figure 7. The source and target subgraphs may belong to different functions or even to different files.
@@
After this initialization, the functions that implement the backward and forward traversals are called (lines 3 and 4). The backwardTraverse() function (lines 7 to 24) iterates over the dboNodes list, which contains the Call nodes related to database operations. For each visited node (dboNode), a Read, Insert, Update, or Delete operation is instantiated, depending on the type of database operation (line 9). The created DatabaseOperation instance (dbo) contains a reference to the corresponding Statement in the Code model, which is obtained from the node.
@@
Next, the Arguments of dboNode are stored in a variable search list sList (line 11), and the control flow graph is traversed backwards from the source node to the current node (lines 12–13). Each visited node (sNode) is processed as follows: If the node corresponds to a database operation call and the variable receiving the result of that call matches one of the variables in sList (line 14), then the current operation (dbo) is connected to the operation created for sNode (pDBO) via the previousDBO relationship (line 16). In turn, pDBO is connected back to dbo through the nextDBO relationship (line 17). This bidirectional connection indicates a data dependency between the two operations—that is, the output of one serves as input to the other. If the two operations act on different collections, the operation receiving the data is marked as a join query (line 18). It is important to note that the findDbOperation function is responsible for retrieving the database operation call statement (Call) referenced by the visited node (sNode). This function navigates from the Control Flow model to the Code model and then checks whether the call is present in the set of database operations stored in the DOS model. If the visited node sNode does not meet the conditions above, and its associated Statement is an assignment, the algorithm checks whether the variable on the left-hand side is in sList (line 19). If so, the variable on the right-hand side of the assignment is added to sList (line 20), thus continuing the tracking of data dependencies through variable propagation.
@@
Once the backward traversal is completed, all database operation Call nodes are revisited using a forward traversal (line 4). In this traversal, only nodes corresponding to Read operations are processed, through the forwardTraverse() function. For each Read node (line 28), a DataStructure is instantiated and aggregated into a Container. Both instances are created if they do not already exist (lines 29 and 30). Additionally, the DataStructure is linked to the corresponding Read operation (line 31). A new search list is also initialized in this traversal (line 32), which is used to collect the data retrieved by the Read operations.
@@
Each subgraph is traversed (lines 34 to 42), starting from the node that follows the database operation node (tNode) (lines 34 and 35). For each visited node, its list of variables is iterated to check which variables are holding values read from the database. To do this, each variable is matched against the elements of sList (line 36). When a match is found, the corresponding Read statement is analyzed to identify the accessed fields of the retrieved object. If a property access is detected, a new Field is created and associated with the current DataStructure (lines 37–39). The type of the Field is determined as follows (line 38):
@@
    If another property is accessed from the current property, this indicates that the current property holds an embedded object, and its type is set to Aggregate.
@@
    If a collection operation is found, the type of the current property is set to Collection.
@@
    Otherwise, the type of the property is set to Attribute —Note that references are not detected during the forward traversal of the graph.
@@
Once the forward traversal is completed, the identification of attributes that are actually references is performed by calling the createReferences function (line 5). This function first collects all Read operations marked as join queries (line 46), that is, those for which the previousDatabaseOperation relationship is not null. This collection is then iterated (lines 47–51), and for each join query, the join condition is analyzed to extract the name of the field involved in the join (lines 48 and 49). A Reference type is then created in the DOS model (line 50). This reference points to the target container and is also associated with the corresponding attribute field previously identified during the forward traversal. Since the corresponding data structures may not yet exist during the initial identification of join queries, reference relationships cannot be extracted at that stage. Finally, the algorithm checks whether multiple data structures contain identical sets of fields. When duplicates are found, only one structure is retained and the others are discarded.
Figure 11 shows the DOS model obtained by applying Algorithm 2 to the running example. The DOSmodel aggregates two Read operations —one for each findOne call node— as well as the movies and users containers. Each Read operation is associated with its corresponding data parameter, and each container is linked to a DataStructure that captures the fields of the objects stored within it. The construction of this DOS model proceeds as follows. During the backward traversal, the two Read operations are identified and connected via the next and previous relationships. Since they satisfy the join condition, the second operation is marked accordingly as part of a join query. In the forward traversal, two Container elements are detected, each associated with its respective DataStructure. These containers are identified by inspecting the arguments of the corresponding Read operations. Field detection proceeds in two steps: (i) the name field in the User container and the _id field in the Movie container are identified from the first argument of the respective Read operations, and (ii) the remaining fields are inferred from the usage of the result variables in the if-then Selection node. These fields appear in the condition expression and its only branch, as well as in the three console.log Calls that form the body of the block (see Figure 7). The watchedMovies field is identified as a collection (array) containing elements of a third data structure. Its type is inferred as an array of Aggregate objects, each containing the fields stars and movie_id. The latter is eventually identified as a reference.
@@
We tested Algorithm 2 using small code snippets structurally similar to the running example. For each snippet, we manually verified that the resulting DOS model correctly represented the Container, DataStructure, Field, and Type elements. We began with a minimal example containing a single Read database operation with one property in the filter. The resulting model was checked to ensure that it included one Read operation and one Container, associated with a DataStructure containing a single Field. Subsequently, we incrementally added statements that manipulated the result object returned by the query. This led to the discovery of additional fields, and we re-examined the model to confirm that these fields were properly captured. The process continued in several iterations by progressively incorporating different types of database operations and accessing up to three distinct containers. At each step, new fields were introduced to increase variability in the data structures and to validate the model’s ability to adapt to structural changes.

### v3

@@ 1 Introduction
As far as we know, no approaches applying code analysis to infer NoSQL complete schemas have been published. However, static code analysis of MongoDB applications has been used for related purposes, such as discovering entity evolution and detecting database access operations [20, 6]. [20] presented a strategy aimed at tracking how the schema of a specific entity type evolves throughout an application’s lifecycle. Their approach analyzes different versions of Java applications to identify structural changes over time —i.e., entity versions. However, it does not extract a complete database schema, as relationships between entities are not identified. Boris Cherry et al [6]. proposed a method to detect MongoDB access operations in JavaScript applications. Their work focuses on locating access points—such as queries, insertions, updates, and deletions—across diverse codebases.
@@
    To the best of our knowledge, this is the first code analysis approach capable of extracting NoSQL logical schemas, where entity types are connected through both aggregation and reference relationships. Additionally, the approach supports relationship types. By representing schemas using a generic metamodel, our method becomes applicable to multi-model data-base tools, enabling broader use in diverse database environments.
@@
    The extracted information in the devised reverse engineering process also enables the automation of data-base refactorings. In this paper, we illustrate this capability with a strategy to duplicate fields in order to eliminate the need for expensive joins, thus improving query performance.
@@
    The proposed reverse engineering process leverages metamodel-based abstractions [4] to represent the involved information: source code, control flow, data-base operations, and data structures. The metamodels were designed to ensure platform independence and reusability. In particular, the code metamodel captures common elements of object-oriented languages and can be extended with language-specific constructs.
@@
The full implementation of our approach, which includes all metamodels, model-to-model transformations, and code analysis algorithms used for generating intermediate models, extracting schemas, and automating refactorings, is publicly available on GitHub  [^7] U-Schema: Code analysis Inference and Refactoring - ModelUM Repository: https://github.com/modelum/uschema-code-analysis.
@@ 2 Overview of the approach
User objects contain a watchedMovies field, which holds an array of embedded objects. Each embedded object includes two fields: movie_id, which stores the identifier of the watched movie (i.e., a reference); and stars, which contains the user’s rating of that movie.
@@
1 user = db("Users").query(name == "Brian")
2 if (user.watchedMovies[0].stars >= 5)
3 println user.name + user.surname + user.email
4 println "Lastwatchedmovie":
5 println user.watchedMovies[0].movie_title
@@
1 const MongoDB = require(’mongodb’).MongoClient;
@@
3 const url = ’mongodb://modelum.es/db:27017’;
4 const dbName = ’streamingservice’;
@@
6 const client = new MongoDB(url);
7 client.connect(err => {
@@ 3 Obtaining an abstract representation of the code
Following the principles of Model-Driven Engineering (MDE) [13, 15, 27], we represent the source code to be analyzed through two complementary metamodels, which are described in this section. The Code metamodel captures the structural elements of the application—such as classes, methods, and statements—while the Control Flow metamodel represents the execution order of these statements. These representations provide the foundation for subsequent analysis stages, including the identification of database operations and the extraction of the schema.
@@ 3.1 From source code to Code models
Building on our prior experience with the Java metamodel of the MoDisco framework and the Code metamodel defined in the OMG’s Knowledge Discovery Metamodel (KDM) specification [10], the Code metamodel was designed to provide an abstract and uniform representation of programs, independent of any specific programming language. The former helped us identify and formalize the main object-oriented constructs, while the latter inspired the design of a language-independent representation.
The metamodel captures the the main concepts and relationships that are common to object-oriented languages —such as Statement, MethodInvocation, and Class— providing a lightweight and language-independent representation well suited for schema inference across different programming languages. An excerpt of the metamodel is shown in Figure 4. The metamodel was not tailored to any specific language. Instead, language-specific elements could be defined in separate extension metamodels, allowing for a modular architecture easily adaptable to different programming languages. This design also ensures that the most common statements are centralized in the core Code model, promoting consistency and reuse across language variants.
Next, we describe the Code metamodel in sufficient detail to explain how Control Flow models are derived. A Code model represents an executable unit, such as a JavaScript script or program. It aggregates Containers, global variable declarations (VariableDec), as well as Types. These can be either primitive types or classes. Each Variable has an associated Type. This type is either taken from the variable declaration or locally inferred from expressions such as variable initializations or literal comparisons. The inference of more complex typing information is part of the analysis process and remains outside the scope of this work. A Container represents a structural unit that holds scripts or classes, such as packages, namespaces, folders, or files. Containers can be nested and may aggregate CodeContainers, which in turn group code blocks, class declarations, and variable declarations. A CodeBlock contains an ordered list of Statements, such as conditionals or loops, and also declares local variables. Figure 4 shows only the types of statements that appear in the running example: conditional selection, method calls, variable accesses, and object creation expressions. A special kind of code block is the CallableBlock, which represents executable blocks invoked by the program, such as methods, functions, constructors, and lambda expressions.
@@
It is worth noting that the injector responsible for obtaining Code models from the AST was manually implemented, rather than automatically generated using a language workbench. This decision was motivated by the complexity and size of the Code metamodel.
@@
To validate this stage, we applied the testing strategy defined in [10]. We executed simple tests on minimal code snippets representing specific constructs (e.g., loops), automatically regenerating code from the corresponding Code model and comparing it to the original using a text-diff tool. This iterative process helped verify the correctness of the model injection logic for each construct.
@@ 3.2 Representing the control flow
Code analysis typically requires not only a representation of the syntax tree, but also knowledge of the control flow graph. To represent control flow in our approach, we have defined the metamodel shown in Figure 6. This metamodel defines the abstractions required to represent control-flow information obtained through the Ullman algorithm [1], following the representation strategy we previously applied to PL/SQL triggers [22]. A Control Flow model is derived from a Code model, with its nodes and edges referencing the corresponding statements in the Code model. Both models serve as input to the code analysis process described in the following section. It is important to note that this representation is not a model of the program’s runtime behavior, but rather a control flow model—that is, a structural abstraction of the code’s possible execution paths. This distinction is particularly relevant in the case of JavaScript, where asynchronous constructs may cause the actual execution order to diverge from the control structure represented in the model.
@@
To generate Control Flow models, we adapted the algorithm described in [1]. In our case, the input is a Code model rather than an abstract syntax tree (AST), and the output is a corresponding Control Flow model.
@@
Figure 9: Database Operation&Structure Metamodel.
It is worth noting that the current algorithm does not fully handle callbacks or asynchronous control flows that are resolved only at runtime, which remain outside the current scope of our static analysis.
@@
We validated this second step by visually verifying that the generated models correctly represented the program’s control flow. For this purpose, models were stored in a Neo4J graph database and explored through the Neo4J Browser [10], which displays query results as navigable graphs. Figure 8 shows an excerpt of the Control Flow model for the running example, illustrating an if-then statement with its alternative paths and sequential console.log calls. These visualizations proved much more readable and intuitive than the raw model, facilitating validation. The mapping from the Control Flow model to Neo4J was straightforward, and the generation code was automatically produced using the same mechanism as for the Code model validation.
@@ 4 Finding Operations and Structure of the database
Code and Control Flow models are analyzed to discover the implicit database schema and to apply database refactorings. In the first step of the analysis, information about CRUD operations and the structure of the manipulated data is captured in an intermediate representation. This representation is defined by the Database Operation and Structure (DOS) metamodel, as shown in Figure 9. In this section, we present the Algorithm 1 designed to obtain a DOS model.
@@
Next, we illustrate how the DOS model is incrementally constructed through Algorithm 1, using the running example introduced in Section 2. Table 1 provides a compact step-by-step trace of this process: each row corresponds to a fragment of the source code, the algorithm step in which it is processed, and the resulting update to the DOS model. This stepwise mapping allows readers to visualize how the algorithm progressively shapes the final model depicted in Figure 11.
# Code fragment Algorithm step (line(s)) DOS model update
1 —– Create DOS model (1) Initialize empty DOS model.
2 findOne(users), findOne(movies) Backward Traverse: DB Operation Creation (9). Create two Read object: users and movies, corresponding to the detected database operations.
3 _id == user.watchedMovies[0] .movie_id Backward Traverse: Dependency Detection (18). Detect join condition between users and movies based on field equality; mark movies as a join node.
4 findOne(users), findOne(movies) Forward Traverse: Container and DataStructure Creation (29–30). Create two Containers: users and movies, generate a corresponding DataStructure for each container, and link them to their respective Read object.
5 user.name, user.surname, user.email Forward Traverse: Field Detection (37–39). Create fields name, surname, and email, and add them to container Users.
6 user.watchedMovies[0] .stars >= 5 Forward Traverse: Field Detection (37–39) with new Collection and Container. Create a Collection and a new Container watchedMovies, including the field stars of type Int.
7 movie._id, movie.title Forward Traverse: Field Detection (37–39). Create id and title fields in container.
8 _id == user.watchedMovies[0] .movie_id Reference Creation (48–50). Create a reference between container users and movies.
Table 1: Stepwise mapping from code fragments of the running example to algorithm steps and corresponding DOS model updates.
@@ 5 $createReferences(dboNodes)$
8 Function backwardTraverse(dboNodesdboNodes):
@@ 14     $sNode\leftarrow getPreviousNode(dboNode)$
15     while $\exists$ sNodesNode do
16        if $isDatabaseOperation(sNode)$ $\land$ $getReturnVariable(sNode)\in sList$ then
@@ 20           $markJoinQuery(pDBO)$
22        else if $isAssignment(sNode)$ $\land$ $getLeftVariable(sNode)\in sList$ then
@@ 27       $sNode\leftarrow getPreviousNode(sNode)$
36 Function forwardTraverse(dboNodesdboNodes):
@@ 44     $tNode\leftarrow getNextNode(dboNode)$
45     while $\exists$ tNodetNode do
@@ 52       $tNode\leftarrow getNextNode(tNode)$
61 Function createReferences(dboNodesdboNodes):
@@ 66     $createReference(sField,tField)$
In the following description, ”line X” refers to the corresponding step in Algorithm 1, and ”row Y” indicates the update of the DOS model shown in Table 1.
@@
Before traversing the control-flow graph, two preliminary operations are performed. In an initial step, a DOSmodel root element is created (line 1; row 1). This step is not associated with any code fragment from the running example shown in Listing 1. Next, each subgraph —representing either a function or a script— is traversed to identify the Call nodes associated with database operations. These nodes are collected in the ordered list dbCallNodes (line 2). Relevant nodes are identified by matching function names against those defined in the data-base management API used in the code. It is important to note that in a Control Flow model, each subgraph is connected to the next in the execution sequence through a CallEdge, as shown in Figure 7. The source and target subgraphs may belong to different functions or even to different files.
@@
After this initialization, the functions that implement the backward and forward traversals are called (lines 3 and 4). The backwardTraverse() function (lines 7 to 24) iterates over the dboNodes list, which contains the Call nodes related to database operations. For each visited node (dboNode), a Read, Insert, Update, or Delete operation is instantiated, depending on the type of database operation (line 9; row 2). In the running example, the created objects correspond to the two findOne() operations. The created DatabaseOperation instance (dbo) contains a reference to the corresponding Statement in the Code model, which is obtained from the node.
@@
Next, the Arguments of dboNode are stored in a variable search list sList (line 10), and the control flow graph is traversed backwards from the source node to the current node (lines 12-24). Each visited node (sNode) is processed as follows: If the node corresponds to a database operation call and the variable receiving the result of that call matches one of the variables in sList (line 14), then the current operation (dbo) is connected to the operation created for sNode (pDBO) via the previousDBO relationship (line 16). In turn, pDBO is connected back to dbo through the nextDBO relationship (line 17). This bidirectional connection indicates a data dependency between the two operations—that is, the output of one serves as input to the other. If the two operations act on different collections, the operation receiving the data is marked as a join query (line 18; row 3). In the running example, this corresponds to the equality expression between an object retrieved in the first findOne() operation and the filter condition of the second findOne(). It is important to note that the findDbOperation function is responsible for retrieving the database operation call statement (Call) referenced by the visited node (sNode). This function navigates from the Control Flow model to the Code model and then checks whether the call is present in the set of database operations stored in the DOS model. If the visited node sNode does not meet the conditions above, and its associated Statement is an assignment, the algorithm checks whether the variable on the left-hand side is in sList (line 19). If so, the variable on the right-hand side of the assignment is added to sList (line 20), thus continuing the tracking of data dependencies through variable propagation.
@@
Once the backward traversal is completed, all database operation Call nodes are revisited using a forward traversal (line 4). In this traversal, only nodes corresponding to Read operations are processed, through the forwardTraverse() function. For each Read node (line 28), a DataStructure is instantiated and aggregated into a Container. Both instances are created if they do not already exist (lines 29-30; row 4). In the running example, this corresponds to the creation of two containers, Users and Movies, each one corresponding to a distinct findOne() invocation. Additionally, the DataStructure is linked to the corresponding Read operation (line 31). A new search list is also initialized in this traversal (line 32), which is used to collect the data retrieved by the Read operations.
@@
Each subgraph is traversed (lines 34 to 42), starting from the node that follows the database operation node (tNode) (lines 34 and 35). For each visited node, its list of variables is iterated to check which variables are holding values read from the database. To do this, each variable is matched against the elements of sList (line 36). When a match is found, the corresponding Read statement is analyzed to identify the accessed fields of the retrieved object. If a property access is detected, a new Field is created and associated with the current DataStructure (rows 5, 6 and 7). The type of the Field is determined as follows (line 38):
@@
    If another property is accessed from the current property, this indicates that the current property holds an embedded object, and its type is set to Aggregate (row 6). In the running example, the field watched Movies contains an aggregate object due to the access to its stars property.
@@
    If a collection operation is found, the type of the current property is set to Collection (row 6). In the running example, the field watchedMovies is recognized as a collection of aggregate objects.
@@
    Otherwise, the type of the property is set to Attribute (rows 5 and 7) — note that references are not detected during the forward traversal of the graph. In the running example, this applies to fields such as name, surname, and email in the Users container, and title in the Movies container. These attributes are detected because the code accesses the properties of the corresponding objects, such as user.name or movie.title.
@@
Once the forward traversal is completed, the identification of attributes that are actually references is performed by calling the createReferences function (line 5). This function first collects all Read operations marked as join queries (line 46), that is, those for which the previousDatabaseOperation relationship is not null. This collection is then iterated (lines 47–51), and for each join query, the join condition is analyzed to extract the name of the field involved in the join (lines 48 and 49). A Reference type is then created in the DOS model (row 8). In the running example, this corresponds to the reference between the Users and Movies containers, established through the join condition. This reference points to the target container and is also associated with the corresponding attribute field previously identified during the forward traversal. Since the corresponding data structures may not yet exist during the initial identification of join queries, reference relationships cannot be extracted at that stage. Finally, the algorithm checks whether multiple data structures contain identical sets of fields. When duplicates are found, only one structure is retained and the others are discarded. Figure 11 shows the DOS model obtained by applying Algorithm 1 to the running example.
@@
Algorithm 1 was tested on small code snippets structurally similar to the running example. For each snippet, we manually verified that the resulting DOS model correctly captured the expected Container, DataStructure, Field, and Type elements. Starting with a minimal example containing a single Read operation, we progressively added statements manipulating the query result to uncover new fields and containers. After each iteration, the model was rechecked to ensure that the structural changes were correctly reflected.